OpenStudy (cenaida):

Factor out the greatest common monomial: 14y4 - 15y3 + 8y2 -2y A. 2(7y4 - 8y3 + 4y2 - y) B. y(14y3 - 15y2 + 8y - 2) C. 2y(7y3 - 8y2 + 4y - 1) D. y4(14 - 15y-1 + 8y-2 - 2y-3)

OpenStudy (raffle_snaffle):

what do you think?

OpenStudy (raffle_snaffle):

The answer isn't D

OpenStudy (raffle_snaffle):

Or A

OpenStudy (cenaida):

C

OpenStudy (raffle_snaffle):

why C? Is there any # that can be multiplied by 2 to obtain 15?

OpenStudy (solomonzelman):

it cannot be C, because the 2 times 8 is not equal to 15 Look over choice C

OpenStudy (solomonzelman):

\[14y^4 - 15y^3 + 8y^2 -2y\]\[y(14y^3 - 15y^2 + 8y^1 -2)\]\[2(14y^3 - 15y^2 + 8y -2)\]

OpenStudy (cenaida):

oooh then its B

OpenStudy (solomonzelman):

Yes.

OpenStudy (cenaida):

:) ty

OpenStudy (cenaida):

True or False: To factor a polynomial using the grouping method, factor out the common terms from the first two terms and then the last two terms in the polynomial.

OpenStudy (solomonzelman):

Eliminate A, C and D because of the second term that they each form. thats the prev question....

OpenStudy (solomonzelman):

You mean? \[14y^4 - 15y^3 + 8y^2 -2y\] \[14y^4=y \times y \times y \times y \times 2 \times 7\]\[-15y^3=y \times y \times y \times 3 \times 5 \times (-1)\]\[8y^2=y \times y \times 2 \times 2 \times 2\]\[-2y=y \times 2 \times (-1)\]

OpenStudy (solomonzelman):

Id really get the above statement.

OpenStudy (cenaida):

is it true??

OpenStudy (solomonzelman):

I don't get the statement, so I really can't say }word problems" is my weak spout. give me a second...

OpenStudy (solomonzelman):

Yes, TRUE.

OpenStudy (solomonzelman):

I read it over, and yes.

OpenStudy (cenaida):

thanks!

OpenStudy (solomonzelman):

Anytime, I always try my best to help, although I am better at doing staff rather than explaining.